what is molar solubity of mn(oh)2(ksp=4.5*10) in a buffer solution containing equql amounts of nh4+ and nh3(kb=1.8*10^-5)

Grade:10

1 Answers

Ambarish Kumar Banerjee
7 years ago
NH3+ h20 gives NH4+ +  Oh-. Kb=[NH4+][Oh-]/[NH3]. now [NH4+]=[NH3]. Kb=[Oh-].=given
Ksp=[Mn2+][Oh-](square). now find [Mn2+] active mass. that is equal to solubility=4.5*10/((1.8*10^-5)^2)
if Ksp=4.5*10 as u say.    
really Ksp of manganese hydeoxide is 45???